OpenStudy (anonymous):

Reduce this algebraic fraction 4x^2-y^2/8x^2+10xy+3y^2*4x^2-9xy-9y^2/2x^2-5xy-3y^2

OpenStudy (anonymous):

add like terms

OpenStudy (anonymous):

then do the equation in order of multiplication, division, addition, and subtraction

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

do you understand or no?

OpenStudy (anonymous):

(4x^2 - y^2) / (x^2 - 6xy + 9y^2) ???

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

here let me recruit lol I have confused myself

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

wait now take what you have and complete polynomial division otherwise know as long division

OpenStudy (anonymous):



OpenStudy (campbell_st):

factor the numerator of the 1st fraction, its the difference of 2 squares \[4x^2 - y^2 = (2x - y)(2x + y)\] the denominator of the 2nd fraction can be factored to \[2x^2 -5xy - 3y^2 = (2x + y)(x - 3y)\] substitute these 2 factored forms and then remove any common factors.

OpenStudy (anonymous):

take it away Campberll I figured I was wrong

OpenStudy (campbell_st):

the numerator of the 2nd fraction can also be factored to \[4x^2 - 9xy - 9y^2 = (4x + 3y)(x - 3y)\] see if you can factor the denominator of the 1st fraction

OpenStudy (campbell_st):

so your problem now looks like \[\frac{(2x - y)(2x + y)}{8x^2 + 10xy + 3y^2} \times \frac{(4x+3y)(x - 3y)}{(2x + y)(x - 3y)}\] you just need to factor the denominator of the 1st fraction then remove common factors for the solution hope it helps

OpenStudy (anonymous):

wouldn't it just be 2x-y/x-3y then?

OpenStudy (campbell_st):

well what did you get when you factored the denominator of the 1st fraction... and I think your solution is incorrect...

OpenStudy (anonymous):

(4x+5+1y^2)

OpenStudy (campbell_st):

no that's not close..

OpenStudy (anonymous):

I have no idea then

OpenStudy (campbell_st):

well 3y^2 is prime so the factors are y and 3y find the factors of 8x^2 that when multiplied by y and 3y give the middle term 10xy